### v1

# Attentive VQ–VAE

Abstract
We present a novel approach to enhance the capabilities of VQ-VAE models through the integration of an Attentive Residual Encoder (AREN) and a Residual Pixel Attention layer. The objective of our research is to improve the performance of VQ-VAE while maintaining practical parameter levels. The AREN encoder is designed to operate effectively at multiple levels, accommodating diverse architectural complexities. The key innovation is the integration of an inter-pixel auto-attention mechanism into the AREN encoder. This approach allows us to efficiently capture and utilize contextual information across latent vectors. Additionally, our models uses additional encoding levels to further enhance the model’s representational power. Our attention layer employs a minimal parameter approach, ensuring that latent vectors are modified only when pertinent information from other pixels is available. Experimental results demonstrate that our proposed modifications lead to significant improvements in data representation and generation, making VQ-VAEs even more suitable for a wide range of applications.
Index Terms—  VQ–VAE, Attention, Face Generation, GANs.
1 Introduction
The field of generative models has seen significant advances in recent years, enabling the creation of high-quality and diverse synthetic data [1, 2, 3, 4, 5, 6, 7, 8]. Among these, the Variational Autoencoder (VAE) has emerged as a robust framework for learning latent representations of data distributions [1]. However, traditional VAEs need help generating rich textures and tend to produce over-smoothed images. Thus, refinement models are used to improve the image’s realness. More recently, the Variational Autoencoder with Vectored Quantization (VQ-VAE) [5] was introduced to address VAE’s limitations, offering a novel approach that combines the strengths of autoencoding with discrete vector quantization [9]. The VQ–VAE architecture provides a unique solution to the problem of capturing fine-grained details in data while maintaining the interpretability of latent representations. However, traditional VQ–VAE faces challenges in modeling complex dependencies and preserving long-range consistency in generated samples. With an extra computational cost, a solution consists of implementing a hierarchical codification [5]. Hence, the latent vectors, ${\bf z}_{i}$ in higher codification levels, are computed with an extended support region in the original image.
Among applications of VQ-VAEs are image denoising, data compression [9], data generation [10, 8], abnormality detection [11], and image/video super-resolution [12], to mention a few. Herein, we introduce a variant of VQ–VAE by incorporating an Attention and Hierarchical mechanisms [13, 14] that extends the codification capabilities. These advancements enable the encoder to more effectively preserve intricate features in the generated samples. For instance, in image generation, Attentive VQ–VAE demonstrates improved capabilities in capturing subtle facial features, like the symmetry of facial attributes, color distribution of eyes, and nuanced contours of facial components. This paper presents the architecture and training strategy of Attentive VQ–VAE, demonstrating its effectiveness through extensive experimental results. Incorporating a Generative Adversarial Network (GAN) training strategy further enriches the model’s capabilities by reducing the required training iterations. Our numerical experiments highlight the distinct advantages of Attentive VQ–VAE over its predecessors.
Fig. 1: General scheme of a VQ–VAE
The remainder of this paper is organized as follows: Section 2 provides a comprehensive overview of VQ-VAE models. Section 3 details the architecture of the Attentive Hierarchical VQ–VAE, including its novel components and improvements over the original VQ–VAE framework. That Section also delves into the training strategy based on GANs techniques [15]; indeed, a PatchGAN [16]. Section 4 presents the experimental setup and showcases the results achieved by investigated versions of Attentive Hierarchical VQ–VAE, quantitatively and qualitatively comparing their performance. Finally, Section 5 concludes the paper by summarizing the contributions and discussing potential future research directions in the Attentive VQ–VAE and generative modeling.
2 Related Work
In Fig. 1, we depict the general scheme of the VQ-VAE. In this, the encoder (ENC) transforms the input data from its original space into an array of latent vectors ${\bf z}$ (in dimension $N$) with minimal information loss. Subsequently, a Vector Quantizer (VQ) replaces each vector ${\bf z}_{i}\in{\bf z}$ with the vector
${\bf e}_{k}=\arg\min_{{\bf e}_{k}\in D}\|{\bf z}_{i}-{\bf e}_{k}\|,$ (1)
where $D$ is a dictionary of vectors learned from the data. Hence, each latent vector can only be one of those defined by the dictionary. In this architecture, such a dictionary is learned from the data at training time. Thus, if the dictionary $D$ exhibits sufficient diversity and the encoder effectively maps ${\bf z}$ to ${\bf e}$ with minimal error, then the decoder (DEC) is capable of generating a reliable reconstruction of the original data. An important distinction between the VQ-VAE and the traditional VAE lies in their treatment of latent variables: the former quantizes them [17, 5], while the latter constraint the latent variables by imposing them a prior distribution, often adopting a multivariate normal distribution with zero mean and identity covariance: $p({\bf z})\sim\mathcal{N}(0,I)$.
VQ–VAE models have been noted for their efficiency in encoding data into low-dimensional latent spaces. Once one trains a VQ–VAE model, the prior distribution of the latent variables, $p({\bf z})$, can be learned using autoregressive models such as PixelCNN [3, 18]. Then, one can generate new data by sampling $p({\bf z})$ and decoding such samples. Our work focuses solely on improving the efficiency and capacity of VQ-VAE models. It is essential to mention that the prior distribution $p({\bf z})$ estimation is beyond our study’s scope. We focus on the VQ–VAE encoder’s inherent loss of information challenge. In the context of facial images, VAE-based models have exhibited deficiencies in global consistency [10, 8]. These issues encompass generated faces with asymmetrical features (e.g., eyes of different colors) and overall incoherence (e.g., disproportional features). One can attribute these challenges to the latent vectors derived from convolutional networks, which capture characteristics from localized regions. However, enlarging the support region implies expanding convolutional kernels, leading to an escalation in parameters and training time.
3 Method
Our focus in this paper centers on enhancing the encoder’s capabilities by redesigning encoders, reducing encoding resolution levels, and integrating an attention mechanism. These modifications aim to augment the performance of the VQ-VAE while maintaining parameters at practical levels.
3.1 Attentive Residual Encoder
Fig. 2 depicts the schematic of our proposed Attentive Residual Encoder (AREN). Our encoder design draws inspiration from the multilevel encoder proposal of VQ-VAEv2. Although Fig. 2 illustrates the two-level case, our implementation accommodates multiple levels (we have tested up to three). The dashed rectangle represents the encoder within the broader scheme, as shown in Fig. 1. Notably, the base encoder’s output is shared among all encoder levels. We design the AREN-type encoders to operate complementary by effectively splitting the information corresponding to each level. The latent vector of the upper level undergoes quantization (VQ2) and scaling (RZ2) to align the Height and Width dimensions with those of the lower level. Then, we concatenate, by channels, the quantized response of the upper level with the AREN response of the lower level. We combine the concatenated channels with a 1x1 convolution. This processed tensor serves as the output of the proposed encoder and is passed to the vector quantizer.
In Panel (a) of Fig. 3, we illustrate the constituents of the AREN encoder. It adopts the convolutional residual network architecture ResNetv2 [19], adding a pixel attention layer and a 1x1 convolution to adjust channel numbers. Particularly noteworthy is the inter-pixel auto-attention layer inspired by PixelAttention; see Panel (b) in Fig. 3.
Additional Encoding Levels. To introduce an extra encoding level, a hypothetical level 0 in Figure 2, we take the output of the current lower level and adjust its dimensions to match the AREN output of the new lower level. Thus, concatenate by channels those tensors. The remaining aspects of the new lower level closely resemble those of the level immediately above it.
Fig. 2: Attentive Residual Encoder (AREN).
Fig. 3: Modules of AREN: (a) Residual Encoder. and (b) Self–attention module.
3.2 Residual Pixel Attention
The attention layer objective is to incorporate information from similar pixels into each latent vector through a residual-based strategy. Our attention layer employs minimal parameters, and the residual approach ensures modifications to the latent vector only when pertinent information from other pixels is available.
Given the tensor $x$, with dimensions $(h,w,f)$, the attention matrix $W$ is computed as
$W_{ij}=\sigma(g_{1}(x_{i})\,g_{2}(x_{j})^{\top}),$ (2)
where $g_{1}$ and $g_{2}$ are convolutional layers with kernel size equal to $1\times 1$, $\sigma$ denotes the sigmoid activation function, and the number of filters equals the latent space dimension, $c$. Then, the residual attention layer implements:
$x\leftarrow x+Wx$ (3)
The pseudocode in Algorithm 1 presents the details of the Attention layer. Such a pseudocode computes self-attention if we pass $y=x$ as parameters; in another case, a cross-attention.
Algorithm 1 PixelAttentionV2
Input Tensor $x$ with dimensions $(b,h,w,f)$
      Input Tensor $y:(b,h,w,f)$
      Output Updated $x:(b,h,w,c)$
1:procedure Attention($x,y$)
2:     $y\leftarrow\textit{Conv2D}(c,1\times 1)(y)$triangleright Num filres $c$
3:     $x\leftarrow\textit{Conv2D}(c,1\times 1)(x)$
[…]
6:     $W_{b,i,j}\leftarrow x_{b,i,c}\,y_{b,j,c}$triangleright Einstein notation
7:     $W\leftarrow\sigma(W)$
8:     $x_{b,j,c}\leftarrow x_{b,j,c}+W_{b,i,j}\,y_{b,j,c}$
[…]
10:return $x$
                   (a) GT                            (b) H-VQ-VAE                  (c) Attentive H-VQ-VAE            (d) Attentive VQ-VAE
Fig. 4: (a) Random CelebA-HQ images $256\times 256$ pixels, and their reconstructions computer with the proposed method: (b) Hierarchical without attention and two encoding levels; (c) Attentive–Hierarchical VQ-VAE with two encoding levels; and (d) Attentive VQ-VAE with one encoding level.
3.3 Implementation Details
Assuming input images of $256\times 256$ pixels with the color channels $(H,W,3)$, Base-Residual-Encoder transforms them to a tensor $(H,W,C)$
We inspire on ResNetV2 to implement our blocks ”Id-ResBlock” and ”Conv-ResBlock”. Let us break down the structure of these blocks:
  1. 1.
    Id-ResBlock (Identity Residual Block). Layers: BatchNormalization, LeakyReLU Activation (with a slope of $\alpha=0.1$), Convolution2D (with the number of filters equals the number of input channels), and the output of the last convolution is summed with the input to the block. Hence, this block’s input and output have matching dimensions.
  2. 2.
    Conv-ResBlock (Convolutional Residual Block). This block consists of the same layers as Id-ResBlock, with the difference that the Convolution2D has a stride equal to (2,2). With this stride, the width and height dimensions are reduced by half. For this reason, a Convolution2D with the same parameters is applied to the input data to be added to the output of the main execution path.
In summary, both ”Id-ResBlock” and ”Conv-ResBlock” are designed to facilitate the flow of information through deep neural networks while addressing the vanishing gradient problem. The ”Id-ResBlock” maintains input and output dimensions, while the ”Conv-ResBlock” reduces dimensions through a stride in the convolutional layer and adjusts the input accordingly to allow for summation. These blocks are crucial for enabling the training of deep neural networks effectively. Table 1 summarizes the reminder parameters; the encoders are of the kind residual with the distinction that the Base-Encoder ended with a $1\times 1$ convolution with many filters equal to the latent dimension instead of an attention module of the ARENs. The Discriminator is a convolutional network with strides indicated in the next row of the table.
Convolutional
Module R,r filters Output
Base Res-Encoder 3,2 (128,128,128), +256 (64,64,266)
AREN 1 2,2 (128,128) (32,32,256)
AREN 2 3,2 (128,128,128) (16,16,256)
AREN 3 3,2 ($,128$) (8,8,256)
Discriminator (128,128,128,64,64,1) (32,32,1)
(strides) (2,2,2,1,1,1)
Table 1: Summary of model hyper-parameters.
4 Experiments
In this work, for demonstration purposes, we focus on face generation using the CelebA-HQ dataset[20] with a resolution equal to $256\times 256$ pixels. We used 80% of the images for training and the remaining 20% for testing.
Fig. 4 depicts generated faces; as we can see, these are very similar to the input faces. Panel (a) shows random examples of the test set (Ground True, GT). The following panels show the reconstructions computed with our proposal Attentive and Hierarchical VQ–VAE:
[…]
We trained the models for 400 epochs. Our Hierarchical VQ-VAE (two-level) preserves the textures slightly better than the version with attention. However, the model with attention achieves better symmetry in the generated faces; we can note it when comparing the colors of both eyes. We noted that the attention mechanism alone was sufficient to incorporate long-range relationships between regions of the images. Therefore, we simplified the model by leaving a single level and the attention module, Fig 4(d). The attentive model preserves better face symmetry.
Table 2 shows the computational resources demanded for each model. We train the models in NVidia 3090 RTX for 400 epochs. The patch-discriminator architecture ($0.412$ Millions of parameters) was the same for all the models. Since all the models were trained for 400 epochs, it is reasonable to expect that the AH-VQVAE model has not reached the same grade of convergence as the simplest model. That could explain some observed asymmetries.
Num. parameters Training time
Model (millions) (secs. per epoch )
H-VQVAE 12.446 923
AH-VQVAE 12.496 1195
A-VQVAE 7.549 815
Table 2: Computational resources.
5 Conclusion
We have introduced the Attentive VQ-VAE to enhance the capabilities of VQ-VAE models. The motivation behind this work arises from the need to address limitations in existing generative models, particularly those related to fine-grained details and global consistency in generated images. Our proposed Attentive VQ-VAE incorporates attention and hierarchical mechanisms. These additions aim to improve the encoding and generation capabilities of VQ-VAEs significantly. Through experiments, we demonstrated the effectiveness of the Attentive VQ-VAE in generating high-quality and realistic face images while simultaneously reducing the computational cost. Our attentive model exhibited similar performance in symmetry, color distribution, and facial feature preservation compared to Hierarchical VQ-VAE models. Furthermore, we used a strategy based on Generative Adversarial Networks (GANs) that contributes to more efficient and effective training. We confirm adding attention does not significantly increase the number of parameters, although it does increase the computational training time. On the other hand, we can obtain, through attention, results similar in quality to those obtained using hierarchical models while keeping the complexity of the model and its training time under control. In summary, our research showcases the potential of the Attentive VQ-VAE as a valuable tool for various applications, especially in image generation. Our model opens up exciting possibilities in computer vision, image processing, and generative art by addressing the challenges of fine-grained details and global consistency. Looking ahead, our work sets the stage for future research directions in generative modeling. Exploring the Attentive VQ-VAE’s potential in other domains and extending its capabilities for even higher-resolution images represent promising avenues for further investigation. We believe this model can significantly contribute to the advancement of generative models and their applications in various creative and practical domains.
Acknowledges. Work supported by CONAHCYT, Mexico (MR Grant CB-A1-43858, AH Scholarship).